OpenStudy (anonymous):

Solve using the rules for signed numbers. ( +34 ) + ( +86) =

OpenStudy (anonymous):

It is 34+86 = ?.....

OpenStudy (anonymous):

No the + are to show that it is a positive number. - means that it is a negative number.

OpenStudy (anonymous):

If we write a number without a minus sign in front of it means that it is positive...

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i don't understand any of it. would you like for me to tell you what my textbook says?

OpenStudy (anonymous):

Yeah sure tell that..:)

OpenStudy (anonymous):

k let me find it...

OpenStudy (anonymous):

When adding two numbers with the same sign, add the numbers like normal, and keep the same sign as the answer. Example: (+2) + (+5) = ( +7 ) Example: (-2) + (-5) = (-7)

OpenStudy (anonymous):

Yeah it's correct..The text book has explained it clearly...:) But normally we don't need to put a + mark in front of a number..

OpenStudy (anonymous):

So how would I add them?

OpenStudy (anonymous):

Just add the two numbers...